OpenStudy (anonymous):

An exam has 12 problems. How many ways can (integer) points be assigned to the problems if the total of the points is 100 and each problem is worth at least 5 points?

OpenStudy (anonymous):

In other words I think they are asking how many ways can the leftover 40 points (since each has to be 5 so 5*12=60) can be assigned to the questions

OpenStudy (anonymous):

I think the answer is C(40+12-1,12-1)=c(51,11) but I don't quite understand why they minus 1

OpenStudy (anonymous):

but can't it also be 51!/40!?

OpenStudy (anonymous):

i think it is what you said

OpenStudy (anonymous):

12 boxes in which to put 40 points

OpenStudy (anonymous):

But how would I do that? do you understand why they subtracted 1? or is that not the correct way to do it

OpenStudy (zarkon):

for 12 boxes you only need 11 dividers

OpenStudy (zarkon):

for example if you had 3 boxes and one item (call it x) was can put it in the 3 boxxes as follows x|| |x| ||x

OpenStudy (anonymous):

that make sense now!! thanks!